# Logarithmic Regret for Online KL-Regularized Reinforcement Learning

### v5

@@ 2 Related Work
Bandits Online Iterative GSHF [^1] Xiong et al. (2024a) studies relative-preference feedback, so their sample complexity additionally depends on $e^{\eta}$. Since our work focuses on absolute rewards, we omit this dependence. (Xiong et al., 2024a) -- $\widetilde{\mathcal{O}}\big{(}d^{2}/{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\epsilon^{2}}\big{)}$$\times$
Two-Stage Mixed-Policy Sampling (Zhao et al., 2024) -- $\widetilde{\mathcal{O}}\big{(}(\eta^{2}C_{\mathrm{GL}}^{2}+\eta/{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\epsilon})\log N_{\mathcal{R}}\big{)}$
@@
MDPs Online Iterative M-GSHF [^2] Xiong et al. (2024b) considers the trajectory-level reward and learns the reward in $\mathcal{R}$ and transition probability in $\mathcal{P}$ separately, so the dependence on $d_{\mathcal{R}},N_{\mathcal{R}}$ and $d_{c}P,N_{\mathcal{P}}$ are separete. (Xiong et al., 2024b) $\widetilde{\mathcal{O}}\big{(}\sqrt{d_{\mathcal{R}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}T}\log N_{\mathcal{R}}}$ $+Hd_{\mathcal{P}}\log N_{\mathcal{P}}\big{)}$ $\widetilde{\mathcal{O}}\big{(}d_{\mathcal{R}}\log N_{\mathcal{R}}/{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\epsilon^{2}}$ $+Hd_{\mathcal{P}}\log N_{\mathcal{P}}/\epsilon\big{)}$$\times$
XPO (Xie et al., 2024) [^3] Xie et al. (2024) considers the deterministic-transition setting and optimizes the policy directly. -- $\widetilde{\mathcal{O}}\big{(}d_{\mathcal{F}}\log N_{\mathcal{F}}/{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\epsilon^{2}}\big{)}$$\times$
@@
For a finite function class $\mathcal{F}$, we use $N_{\mathcal{F}}$ to represent its cardinality. We use $\widetilde{\mathcal{O}}$ to omit the logarithmic orders. We use the convention $[n]=\{1,\ldots,n\}$. For any vector $x$ and a matrix $\Sigma$, let $\|x\|_{\Sigma}=\sqrt{x^{\top}\Sigma x}$. For a function $R:\mathcal{X}\times\mathcal{A}\rightarrow\mathbb{R}$, parameter $\eta>0$ and the reference policy $\pi_{\mathrm{ref}}$, let the normalization constant $Z_{R}(x)=\mathbb{E}_{a\sim\pi_{\mathrm{ref}}(\cdot|x)}\exp(\eta R(x,a))$.
@@ 3.1 KL-Regularized Contextual Bandits
The uncertainty $U_{\mathcal{R}}(\lambda,x,a;\mathcal{D}_{t})$ measures how much difference the new-coming data $x,a$ has with the history data $\mathcal{D}_{t}$, and is widely adopted in RL literature with general function approximation (Zhang, 2023; Ye et al., 2023; Agarwal et al., 2023; Zhao et al., 2023a). To illustrate it better, we use the linear function as a special case, where the function class $\mathcal{R}$ can be embedded into a linear mapping =$\mathcal{R}=\{\theta^{\top}\phi(\cdot,\cdot):~{}\theta\in\mathbb{R}^{d},~{}\|\theta\|_{2}\leq B\}$. Let the covariance matrix $\Sigma_{t}=\sum_{i=1}^{t}\phi(x_{i},a_{i})\phi(x_{i},a_{i})^{\top}+\lambda/B\cdot I$. Then, the eluder coefficient can be simplified as
@@
$J(\pi):=$ $\mathbb{E}_{x\sim d_{0}}\mathbb{E}_{a\sim\pi(\cdot|x)}\Big{[}R^{*}(x,a)-\frac{1}{\eta}\log\frac{\pi(a|x)}{\pi_{\mathrm{ref}}(a|x)}\Big{]}$
@@ 3.2 KL-Regularized Reinforcement Learning
In this section, we introduce the KL-regularized MDP problem. A Markov Decision Process (MDP) is defined by a tuple $\mathcal{M}=({\mathcal{S}},\mathcal{A},H,\mathbb{P},d_{0},r)$, where ${\mathcal{S}}$ is the state space, $\mathcal{A}$ is the action space, $H$ is the time horizon, transition probability =$\mathbb{P}=\{\mathbb{P}_{h}\}_{h=1}^{H}$ denotes the probability $\mathbb{P}(s_{h+1}|s_{h},a_{h})$ of transition from the current $(s_{h},a_{h})$ to the next state $s_{h+1}$ at each step $h$, $d_{0}$ is the initial state distribution, and $r=\{r_{h}:{\mathcal{S}}\times\mathcal{A}\rightarrow\mathbb{R}\}_{h=1}^{H}$ is the reward function.
@@
$\qquad-\frac{1}{\eta}\cdot\mathrm{KL}\bigl{(}\pi_{h^{\prime}}(\cdot|s_{h^{\prime}})\|\pi_{\mathrm{ref},h^{\prime}}(\cdot|s_{h^{\prime}})\bigr{)}\,\Big{|}\,s_{h}\Big{]},$
@@
$\qquad-\frac{1}{\eta}\cdot\mathrm{KL}\bigl{(}\pi_{h^{\prime}}(\cdot|s_{h^{\prime}})\|\pi_{\mathrm{ref},h^{\prime}}(\cdot|s_{h^{\prime}})\bigr{)}\,\Big{|}\,s_{h},a_{h}\Big{]}.$
@@
$\qquad\qquad-\frac{1}{\eta}\cdot\mathrm{KL}\bigl{(}\pi_{h}(\cdot|s_{h})\|\pi_{\mathrm{ref},h}(\cdot|s_{h})\bigr{)}\Big{]},$
@@
${\mathcal{T}}_{\eta}^{h}f_{h+1}(s_{h},a_{h}):=r_{h}(s_{h},a_{h})+\frac{1}{\eta}\mathbb{E}_{s_{h+1}|s_{h},a_{h}}V^{f}_{h+1}(s_{h+1}),$
@@ 4.1 The Proposed Algorithm and Result
$\qquad+\underbrace{\widehat{R}_{t}(x,\pi_{t})-R^{*}(x,\pi_{t})+b_{t}(x,\pi_{t})}_{\displaystyle{\leq 2b_{t}(x,\pi_{t})}}$
$\qquad+\big{(}\widehat{R}_{t}(x,\pi^{*})+b_{t}(x,\pi^{*})-\mathrm{KL}(\pi^{*}(\cdot|x)\|\pi_{\mathrm{ref}}(\cdot|x))\big{)}$
$\qquad-\big{(}\widehat{R}_{t}(x,\pi_{t})+b_{t}(x,\pi^{*})-\mathrm{KL}(\pi_{t}(\cdot|x)\|\pi_{\mathrm{ref}}(\cdot|x))\big{)},$
where on the right-hand side of the last equality, the first two terms follow from the optimistic reward estimator and the definition of the bonus, the subtraction of the last two terms is upper bounded by 00 since $\pi_{t}$ is the maximizer of the KL-regularized reward according to Algorithm 1. Essentially, deserting the last two terms is where they loosen and reduce to the traditional analysis. As a result, their regret is upper bounded by the first-order summation of bonuses $2\sum_{t=1}^{T}\mathbb{E}_{x\sim d_{0}}b_{t}(x,\pi_{t})$ and finally get an $\widetilde{\mathcal{O}}(\sqrt{T})$ regret.
@@
$\quad-\mathbb{E}_{\pi_{t}}\bigg{[}R^{*}(x,a)-\frac{1}{\eta}\log\frac{\pi_{t}(a|x)}{\pi_{\mathrm{ref}}(a|x)}\bigg{]}$
@@
$\quad+\mathbb{E}_{\pi_{t}}\big{[}\widehat{R}_{t-1}(x,a)-R^{*}(x,a)+b_{t-1}(x,a)\big{]},$
@@
$\Delta(x,R):=-\frac{1}{\eta}\log Z_{R}(x)+\mathbb{E}_{\pi_{R}^{\eta}}\big{[}R(x,a)-R^{*}(x,a)\big{]}.$ (5)
@@
$\qquad-\eta\sum_{a^{\prime}\in\mathcal{A}}\pi_{f_{\lambda}}^{\eta}(a|x)\cdot\pi_{f_{\lambda}}^{\eta}(a^{\prime}|x)\cdot\big{(}f_{\lambda}(x,a^{\prime})-R^{*}(x,a^{\prime})\big{)}\Big{)}$
$\qquad\qquad\cdot\big{(}\widehat{R}_{t-1}(x,a)+b_{t-1}(x,a)-R^{*}(x,a)\big{)}\Big{]},$
@@
Under Algorithm 1 and Assumption 3.2 and the condition that the noises $\epsilon_{t}$ are conditional $1$-sub-Gaussian, we have with probability at least $1-\delta$ for all $t\in[T]$, the uniform optimism event that $\mathcal{E}_{\mathrm{opt}}^{t}=\{\widehat{R}_{t}(x,a)+b_{t}(x,a)-R^{*}(x,a)\geq 0,~{}\forall(x,a)\in\mathcal{X}\times\mathcal{A}\}$ holds true.
@@
$\quad\cdot\underbrace{\bigl{(}f_{\lambda}(x,a)-R^{*}(x,a)\bigr{)}\cdot\pi_{f_{\lambda}}^{\eta}(a|x)}_{U(x,a)}\big{]}$
@@
$\qquad\bigl{(}\widehat{R}_{t-1}(x,a)+b_{t-1}(x,a)-R^{*}(x,a)\bigr{)}^{2}\big{]},$
@@
$\qquad\qquad\bigl{(}\widehat{R}_{t-1}(x_{t},a_{t})+b_{t-1}(x_{t},a_{t})-R^{*}(x_{t},a_{t})\bigr{)}^{2}\big{]}$
@@ 5.1 The Proposed Algorithm and Result
$\qquad\qquad\sum_{i=1}^{t-1}(f_{h}(s_{h}^{i},a_{h}^{i})-\widehat{f}_{h}^{t}(s_{h}^{i},a_{h}^{i}))^{2}+\lambda\leq(\beta_{h}^{T})^{2}\Big{\}},$
@@
$\qquad+{\mathcal{T}}^{1}\widehat{V}_{2}(s_{1},a_{1})-{\mathcal{T}}^{1}V_{2}^{\widehat{\pi}}(s_{1},a_{1})\big{]}$
@@
$J(\pi^{*})-J(\widehat{\pi})$ $=\sum_{h=0}^{H-1}\mathbb{E}_{s_{1}\sim d_{0}}\underbrace{\bigl{[}V_{1}^{\widehat{\pi}^{(h)}}(s_{1})-V_{1}^{\widehat{\pi}^{(h+1)}}(s_{1})\bigr{]}}_{I_{h+1}}.$
@@
$\qquad-\bigl{[}Q_{h+1}^{\pi^{*}}(s_{h+1},\widehat{\pi}_{h+1})-\eta^{-1}\mathrm{KL}(\widehat{\pi}_{h+1},\pi_{\mathrm{ref},h+1})\bigr{]}\Big{\}}$
@@ Appendix A Proofs for KL-Regularized Contextual Bandits
$\Delta(x,R):=$ $-\frac{1}{\eta}\log Z_{R}(x)+\sum_{a\in\mathcal{A}}\frac{\pi_{\mathrm{ref}}(a|x)\cdot\exp(\eta R(x,a))}{Z_{R}(x)}\cdot\bigl{(}R(x,a)-R^{*}(x,a)\bigr{)}$ (12)
@@
Under Algorithm 1 and Assumption 3.2 and the condition that the noises $\epsilon_{t}$ are conditional $1$-sub-Gaussian, we have with probability at least $1-\delta$ for all $t\in[T]$, the uniform optimism event that $\mathcal{E}_{\mathrm{opt}}^{t}=\{\widehat{R}_{t}(x,a)+b_{t}(x,a)-R^{*}(x,a)\geq 0,~{}\forall(x,a)\in\mathcal{X}\times\mathcal{A}\}$ holds true.

### v6

@@ 2 Related Work
Bandits Online Iterative GSHF [^1] Xiong et al. (2024a) studies relative-preference feedback, so their sample complexity additionally depends on $e^{\eta}$. Since our work focuses on absolute rewards, we omit this dependence. (Xiong et al., 2024a) - $\widetilde{\mathcal{O}}\big(d^{2}/{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\epsilon^{2}}\big)$ $\times$
Two-Stage Mixed-Policy Sampling (Zhao et al., 2024) - $\widetilde{\mathcal{O}}\big((\eta^{2}C_{\mathrm{GL}}^{2}+\eta/{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\epsilon})\log N_{\mathcal{R}}\big)$
@@
MDPs Online Iterative M-GSHF [^2] Xiong et al. (2024b) considers the trajectory-level reward and learns the reward in $\mathcal{R}$ and transition probability in $\mathcal{P}$ separately, so the dependence on $d_{\mathcal{R}},N_{\mathcal{R}}$ and $d_{c}P,N_{\mathcal{P}}$ are separete. (Xiong et al., 2024b) $\widetilde{\mathcal{O}}\big(\sqrt{d_{\mathcal{R}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}T}\log N_{\mathcal{R}}}$ $Hd_{\mathcal{P}}\log N_{\mathcal{P}}\big)$ $\widetilde{\mathcal{O}}\big(d_{\mathcal{R}}\log N_{\mathcal{R}}/{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\epsilon^{2}}$ $Hd_{\mathcal{P}}\log N_{\mathcal{P}}/\epsilon\big)$ $\times$
XPO (Xie et al., 2024) [^3] Xie et al. (2024) considers the deterministic-transition setting and optimizes the policy directly. - $\widetilde{\mathcal{O}}\big(d_{\mathcal{F}}\log N_{\mathcal{F}}/{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\epsilon^{2}}\big)$ $\times$
@@
For a finite function class $\mathcal{F}$, we use $N_{\mathcal{F}}$ to represent its cardinality. We use $\widetilde{\mathcal{O}}$ to omit the logarithmic orders. We use the convention $n]=\{1,\ldots,n\}$. For any vector $x$ and a matrix $\Sigma$, let $\|x\|_{\Sigma}=\sqrt{x^{\top}\Sigma x}$. For a function $R:\mathcal{X}\times\mathcal{A}\rightarrow\mathbb{R}$, parameter $\eta>0$ and the reference policy $\pi_{\mathrm{ref}}$, let the normalization constant $Z_{R}(x)=\mathbb{E}_{a\sim\pi_{\mathrm{ref}}(\cdot|x)}\exp(\eta R(x,a))$.
@@ 3.1 KL-Regularized Contextual Bandits
The uncertainty $U_{\mathcal{R}}(\lambda,x,a;\mathcal{D}_{t})$ measures how much difference the new-coming data $x,a$ has with the history data $\mathcal{D}_{t}$, and is widely adopted in RL literature with general function approximation (Zhang, 2023; Ye et al., 2023; Agarwal et al., 2023; Zhao et al., 2023a). To illustrate it better, we use the linear function as a special case, where the function class $\mathcal{R}$ can be embedded into a linear mapping $\mathcal{R}=\{\theta^{\top}\phi(\cdot,\cdot):~\theta\in\mathbb{R}^{d},~\|\theta\|_{2}\leq B\}$. Let the covariance matrix $\Sigma_{t}=\sum_{i=1}^{t}\phi(x_{i},a_{i})\phi(x_{i},a_{i})^{\top}+\lambda/B\cdot I$. Then, the eluder coefficient can be simplified as
@@
$J(\pi)$= $\mathbb{E}_{x\sim d_{0}}\mathbb{E}_{a\sim\pi(\cdot|x)}\Big[R^{*}(x,a)-\frac{1}{\eta}\log\frac{\pi(a|x)}{\pi_{\mathrm{ref}}(a|x)}\Big]$
@@ 3.2 KL-Regularized Reinforcement Learning
In this section, we introduce the KL-regularized MDP problem. A Markov Decision Process (MDP) is defined by a tuple $\mathcal{M}=({\mathcal{S}},\mathcal{A},H,\mathbb{P},d_{0},r)$, where $\mathcal{S}$} is the state space, $\mathcal{A}$ is the action space, $H$ is the time horizon, transition probability $\mathbb{P}=\{\mathbb{P}_{h}\}_{h=1}^{H}$ denotes the probability $\mathbb{P}(s_{h+1}|s_{h},a_{h})$ of transition from the current $(s_{h},a_{h})$ to the next state $s_{h+1}$ at each step $h$, $d_{0}$ is the initial state distribution, and $r=\{r_{h}:{\mathcal{S}}\times\mathcal{A}\rightarrow\mathbb{R}\}_{h=1}^{H}$ is the reward function.
@@
$\hskip 18.49988pt-\frac{1}{\eta}\cdot\mathrm{KL}\bigl(\pi_{h^{\prime}}(\cdot|s_{h^{\prime}})\|\pi_{\mathrm{ref},h^{\prime}}(\cdot|s_{h^{\prime}})\bigr)\,\Big|\,s_{h}\Big],$
@@
$\hskip 18.49988pt-\frac{1}{\eta}\cdot\mathrm{KL}\bigl(\pi_{h^{\prime}}(\cdot|s_{h^{\prime}})\|\pi_{\mathrm{ref},h^{\prime}}(\cdot|s_{h^{\prime}})\bigr)\,\Big|\,s_{h},a_{h}\Big].$
@@
$\hskip 18.49988pt\hskip 18.49988pt-\frac{1}{\eta}\cdot\mathrm{KL}\bigl(\pi_{h}(\cdot|s_{h})\|\pi_{\mathrm{ref},h}(\cdot|s_{h})\bigr)\Big],$
@@
${\mathcal{T}}_{\eta}^{h}f_{h+1}(s_{h},a_{h})=r_{h}(s_{h},a_{h})+\frac{1}{\eta}\mathbb{E}_{s_{h+1}|s_{h},a_{h}}V^{f}_{h+1}(s_{h+1}),$
@@ 4.1 The Proposed Algorithm and Result
$\hskip 18.49988pt+\underbrace{\widehat{R}_{t}(x,\pi_{t})-R^{*}(x,\pi_{t})+b_{t}(x,\pi_{t})}_{\displaystyle{\leq 2b_{t}(x,\pi_{t})}}$
$\hskip 18.49988pt+\big(\widehat{R}_{t}(x,\pi^{*})+b_{t}(x,\pi^{*})-\mathrm{KL}(\pi^{*}(\cdot|x)\|\pi_{\mathrm{ref}}(\cdot|x))\big)$
$\hskip 18.49988pt-\big(\widehat{R}_{t}(x,\pi_{t})+b_{t}(x,\pi^{*})-\mathrm{KL}(\pi_{t}(\cdot|x)\|\pi_{\mathrm{ref}}(\cdot|x))\big),$
where on the right-hand side of the last equality, the first two terms follow from the optimistic reward estimator and the definition of the bonus, the subtraction of the last two terms is upper bounded by 0 since $\pi_{t}$ is the maximizer of the KL-regularized reward according to Algorithm 1. Essentially, deserting the last two terms is where they loosen and reduce to the traditional analysis. As a result, their regret is upper bounded by the first-order summation of bonuses $2\sum_{t=1}^{T}\mathbb{E}_{x\sim d_{0}}b_{t}(x,\pi_{t})$ and finally get an $\widetilde{\mathcal{O}}(\sqrt{T})$ regret.
@@
$\qquad-\mathbb{E}_{\pi_{t}}\bigg[R^{*}(x,a)-\frac{1}{\eta}\log\frac{\pi_{t}(a|x)}{\pi_{\mathrm{ref}}(a|x)}\bigg]$
@@
$\qquad+\mathbb{E}_{\pi_{t}}\big[\widehat{R}_{t-1}(x,a)-R^{*}(x,a)+b_{t-1}(x,a)\big],$
@@
$\Delta(x,R)=-\frac{1}{\eta}\log Z_{R}(x)+\mathbb{E}_{\pi_{R}^{\eta}}\big[R(x,a)-R^{*}(x,a)\big].$ (5)
@@
$\hskip 18.49988pt-\eta\sum_{a^{\prime}\in\mathcal{A}}\pi_{f_{\lambda}}^{\eta}(a|x)\cdot\pi_{f_{\lambda}}^{\eta}(a^{\prime}|x)\cdot\big(f_{\lambda}(x,a^{\prime})-R^{*}(x,a^{\prime})\big)\Big)$
$\hskip 18.49988pt\hskip 18.49988pt\cdot\big(\widehat{R}_{t-1}(x,a)+b_{t-1}(x,a)-R^{*}(x,a)\big)\Big],$
@@
Under Algorithm 1 and Assumption 3.2 and the condition that the noises $\epsilon_{t}$ are conditional 11-sub-Gaussian, we have with probability at least $1-\delta$ for all $t\in[T]$, the uniform optimism event that $\mathcal{E}_{\mathrm{opt}}^{t}=\{\widehat{R}_{t}(x,a)+b_{t}(x,a)-R^{*}(x,a)\geq 0,~\forall(x,a)\in\mathcal{X}\times\mathcal{A}\}$ holds true.
@@
$\qquad\cdot\underbrace{\bigl(f_{\lambda}(x,a)-R^{*}(x,a)\bigr)\cdot\pi_{f_{\lambda}}^{\eta}(a|x)}_{U(x,a)}\big]$
@@
$\hskip 18.49988pt\bigl(\widehat{R}_{t-1}(x,a)+b_{t-1}(x,a)-R^{*}(x,a)\bigr)^{2}\big],$
@@
$\hskip 18.49988pt\hskip 18.49988pt\bigl(\widehat{R}_{t-1}(x_{t},a_{t})+b_{t-1}(x_{t},a_{t})-R^{*}(x_{t},a_{t})\bigr)^{2}\big]$
@@ 5.1 The Proposed Algorithm and Result
$\hskip 18.49988pt\hskip 18.49988pt\sum_{i=1}^{t-1}(f_{h}(s_{h}^{i},a_{h}^{i})-\widehat{f}_{h}^{t}(s_{h}^{i},a_{h}^{i}))^{2}+\lambda\leq(\beta_{h}^{T})^{2}\Big\},$
@@
$\hskip 18.49988pt+{\mathcal{T}}^{1}\widehat{V}_{2}(s_{1},a_{1})-{\mathcal{T}}^{1}V_{2}^{\widehat{\pi}}(s_{1},a_{1})\big]$
@@
[…]
@@
$\hskip 17.00024pt-\bigl[Q_{h+1}^{\pi^{*}}(s_{h+1},\widehat{\pi}_{h+1})-\eta^{-1}\mathrm{KL}(\widehat{\pi}_{h+1},\pi_{\mathrm{ref},h+1})\bigr]\Big\}$
@@ Appendix A Proofs for KL-Regularized Contextual Bandits
$\Delta(x,R)$= $-\frac{1}{\eta}\log Z_{R}(x)+\sum_{a\in\mathcal{A}}\frac{\pi_{\mathrm{ref}}(a|x)\cdot\exp(\eta R(x,a))}{Z_{R}(x)}\cdot\bigl(R(x,a)-R^{*}(x,a)\bigr)$ (12)
@@
Under Algorithm 1 and Assumption 3.2 and the condition that the noises $\epsilon_{t}$ are conditional 11-sub-Gaussian, we have with probability at least $1-\delta$ for all $t\in[T]$, the uniform optimism event that $\mathcal{E}_{\mathrm{opt}}^{t}=\{\widehat{R}_{t}(x,a)+b_{t}(x,a)-R^{*}(x,a)\geq 0,~\forall(x,a)\in\mathcal{X}\times\mathcal{A}\}$ holds true.